Grade 12Algebra

Please help me sir !!!!!!
Sum of the series of 2/2! + 12/4! + 120/6! + 1680/8! +. .............infinity is
1) 2e + 1
2) e-1
3) 2e - 1
4) e + 1

Ranjeet Gupta
7 Years agoGrade 12

1 Answer

Aditya Gupta
ApprovedApproved Tutor Answer7 Years ago
ok so first off it is important to note that 2/2!= 1= 1/1!, 12/4!= 12/24= ½= ½!, 120/6!= 120/720= 1/6= 1/3!, 1680/8!= 1680/40320= 1/24= ¼! …. and hands on.
so the series basically is nothing but 1/1! + ½! + 1/3! + ¼! + …..
= (1 + 1/1! + ½! + 1/3! + ¼! + …..) – 1
= (1/0! + 1/1! + ½! + 1/3! + ¼! + …..) – 1
e – 1
kindly approve :)